# Robust Plan Evaluation based on Approximate Probabilistic Machine Learning

### v2

@@ 5.1.4. Baselines
We compare Roq against a classical RDBMS optimizer, that of IBM Db2 , as well as two prominent recent works, Neo (Marcus et al., 2019) and Bao (Marcus et al., 2024). Neo and Bao were selected due to their demonstrated robustness to errors in the input features. Each of the approaches is evaluated for its ability to select a robust plan from the same set of plans for a given query. To this end, the search space is identical for all approaches and is enumerated as explained.
@@
Neo uses a learned cost model that takes query and plan attributes as input. It feeds the query attributes to a stack of MLP layers and appends the resulting vector to the attributes of every node in the vectorized plan tree. It then feeds the extended plan trees to a stack of TCNN layers followed by another stack of MLPs which gradually reduce the dimensionality of the vector and ultimately predict the execution time.
@@ 5.1.5. Evaluation Measures
We design five sets of experiments to evaluate Roq based on prediction error, correlations with runtime, robustness to workload shifts, the inference overheads, and an ablation study on the role of plan and estimation risks. Our comparisons employ the following measures.
@@
Correlation. We measure the correlations between the estimates (or the optimizer’s Cost) and the actual execution times using Spearman’s rank correlation coefficient rather than the more widely used Pearson’s correlation coefficient. While Pearson’s coefficient measures the linear relationships between two variables, Spearman’s coefficient measures the monotonic relationships, whether linear or not. Therefore, given the estimated costs or execution times are used to rank the plans, Spearman’s coefficient is a more suitable measure of correlation in this context.
Plan Suboptimality. The quality of plans generated by a strategy is evaluated using the Suboptimality measure. The suboptimality of a plan $p_{i}$ is computed with respect to the best performing plan $p*$: $\text{Subopt}(p_{i})=\frac{ET(p_{i})}{ET(p^{*})}\in\mathbb{R};1<\text{Subopt}(p_{i})<\infty$, where $ET(p_{i})$ is the execution time of the plan $p_{i}$, and $p*=argmin_{i\in\{1,\ldots,S\}}ET(p_{i})$. We assume that the best performing plan among the ones generated by the hint sets is $p^{*}$. The distribution of this measure over all plans selected by a certain method, particularly the tail-end of the distribution, characterizes the worst-case performance degradation of that method compared with an oracle which always selects the best plan. This is suggested as a measure of end-to-end robustness (Haritsa, 2020).
Runtime. Improving suboptimality as a measure of robustness may lead to picking plans that are more robust but more expensive than the optimal plan. While the main goal of our work is to improve robustness, we need to demonstrate that robust query optimization does not cause a major slowdown for a given workload. Therefore, the total runtime of the workload is measured for each approach.
Inference Overheads. Using the risk-aware plan evaluation algorithms has an overhead for the inference time since multiple predictions must be made using the base model. This overhead is most significant for plan selection by suboptimality risk given the additional computations needed. Therefore, we measure the inference time for this algorithm for various inference iterations while evaluating the impact on plan suboptimality and runtime.
@@ 5.2. Experimental Results
Prediction Accuracy. The average predictive performance of Roq’s base model (without uncertainty quantification) consistently outperforms the baselines. Roq’s prediction’s q-error is superior to the baselines in all tested scenarios based on CEB, JOB, and TPC-DS, as shown in Figure 4a. Furthermore, the correlation of the predicted runtime with the actual runtime is more significant for Roq compared with either the baseline ML-based approaches or the Db2 optimizer’s cost, as shown in Figure 4b. In these evaluations we only use the raw predictions from each model and do not use MC dropout variational inference to obtain a more accurate prediction.
Figure 4. Roq’s predictive performance vs. the baselines
@@
Model Architecture. Better average predictive accuracy may not necessarily lead to better plan selection performance. Therefore, we perform extensive experiments to evaluate the quality of the generated plans based on their suboptimality and runtime. The plans selected by the risk-aware strategies using Roq demonstrate greater robustness based on plan suboptimality and increased runtime improvements compared with the baselines, as shown in Figure 5.To isolate the impact of the plan selection strategies as implemented by Algorithms 2, 1, and 3, Roq is once tested without risk quantification, similar to the baselines. This scenario is denoted as the base model. On average, Roq’s base model improves the mean and the $99^{th}$ percentile of suboptimality compared with the optimizer by 19.9% and 55.6% respectively. In addition, it improves the same measures compared with the best ML-based baseline by up to 27% and 30% respectively, and it improved runtime compared with the best ML-based baseline by 1.5%. These improvements are attributed to Roq’s architecture using GNNs for producing node embeddings from the join graph, which enables enhanced generalization to the unseen test queries.
@@
Figure 5. The performance of the plans selected by each approach illustrated by (a) suboptimality distribution and (b) runtime improvements, regressions, and net change (all in percentage) for CEB, JOB, and TPC-DS queries
@@
Plan Selection Strategies In comparison with the ML-based baselines, conservative plan selection (denoted by conservative) yields the greatest level of robustness by improving the mean and the $99^{th}$ percentile of suboptimality on average by 39.3% and 35.6% respectively. Plan selection by suboptimality risk (denoted by risk) improves the same measures by 39.4% and 33.7% respectively. These two strategies improve runtime compared with the best ML-based baseline on average by 5.8% and 16.9% respectively. These two strategies are combined with plan pruning based on plan and estimation risks. The pruning thresholds are tuned twice: once by minimizing the total runtime and another time by minimizing the average suboptimality on the validation set, denoted by pruned RN and pruned SO respectively. Neither of the two flavors of pruning has a significant impact on either suboptimality or runtime of the Conservative strategy over the three benchmarks. However, pruning achieved an additional 3% to 5% runtime improvement when used together with the Risk strategy. This is expected as the Conservative strategy adjusts predictions by their uncertainties and effectively eliminates risky predictions by default. However, the Risk strategy benefits more from pruning highly uncertain predictions as computing the SOR for a target plan compared against highly uncertain plans may skew the results and reduce the effectiveness of the predictions.
Using pruning where the uncertainty threshold is obtained by minimizing runtime on the validation set degrades the suboptimality measure, while it reduces runtime regressions (5a). The risk thresholds for pruning plans are tuned such that plans with decent robustness are removed in favor of less robust near-optimal plans. Pruning by minimizing suboptimality on the validation set, however, improves the suboptimality measure both at average and at the tail-end of the distribution. Thus, the choice of pruning strategy comes down to prioritization of robustness over optimality.
@@
Figure 6. Comparing the impacts of the risk-based plan selection strategies using model, data, and total uncertainties in (a) suboptimaility distribution and (b) runtime improvements, regressions, and net change (all in percentage) for TPC-DS queries
@@
Plan vs. Estimation Risks To evaluate the roles of the plan risk (quantified by data uncertainty) and estimation risks (quantified by model uncertainty) on plan selection, we run an ablation study in which the risk-aware plan selection strategies are tested using either model, data, or total uncertainties. We use the extended TPC-DS benchmark for which we observed a more significant impact from the risk-aware plan selection methods. The quality of the resulting plans is evaluated in terms of runtime, improvements or regressions as well as the suboptimality of the resulting plans (Figures 6a and 6b). Using either model or data uncertainty for quantifying risks provides significant improvements compared with the base model and using total uncertainties provides maximum improvements. This suggests that the plan and estimation risks each play a role in improving Roq’s performance and the impact is maximized if they are used together. In addition, the impact of accounting for plan risk is more significant for achieving better runtime performance since it effectively avoids long-running plans and thus significantly reduces regressions.
@@
Robustness to Workload Shifts A major challenge of using ML-based techniques in query optimization is dealing with the problem of generalization to out-of-distribution test samples. A learned model becomes obsolete if the characteristics of the test query differ significantly from the ones used in training. While model retraining is possible, it takes time, and therefore the optimizer would typically need to still rely on the obsolete model in the meantime. We evaluate the impact of a minor and a major workload shift for Roq in comparison to the baselines using two simulations. A minor workload shift is simulated using TPC-DS. The baseline model is trained and tested on 4-join queries, while the model representing a workload shift is trained with 1-,2-, and 3-join queries and tested with 4-join queries. A major workload shift is simulated using the CEB and JOB. The baseline model is trained and tested using the JOB workload, while the model representing a workload shift is trained using the CEB workload and tested using the JOB workload. This constitutes a major workload shift as the query templates in CEB are entirely different than those in JOB. In addition, our CEB queries have up to 12 joins while JOB has up to 28 joins.
The results shown in Figure 7 demonstrate Roq’s robustness to out-of-distribution data caused by workload shift in both situations. Figures 7a and 7b show that while the baseline models’ q-error and correlation degrade significantly with a workload shift, Roq’s predictive performance degradation is minimal. These results confirm a greater robustness of Roq to out-of-distribution data caused by workload shifts. We also evaluate the impact of workload shift on runtime in the major workload shift scenario. As demonstrated in Figure 8a and 8b, the performance of all three models degraded significantly to sub-optimal levels. The Conservative and the Risk strategies, however, significantly reduced regressions leading to performances still superior to the optimizer.
@@
Figure 8. JOB runtime changes after a severe workload shift demonstrated by (a) runtime improvements vs. regressions (b) workload accumulative runtime
[…]
Inference Overheads Capturing estimation risk using the MC Dropout method requires multiple model invocations. A higher number of inferences should provide a more accurate estimate of the expected values and the uncertainties. However, this incurs additional overhead for the query compilation phase which is on the critical path of query execution and must be kept at a minimum. To this end, we run experiments with inference iterations ranging from 5 to 100. We use the suboptimality risk, expected to have a larger overhead compared to the conservative plan selection, given the complex required computations. We capture the inference time as well as the runtime and suboptimality of selected plans. We report an average of 10 runs for each experiment. The inference times are captured using a single thread on the CPU, typically available for query compilation. The inference times grow linearly with the number of iterations (Figure 9a). Figures 9b and 9c illustrate the impact of increasing the number of inference iterations on the suboptimality and the runtime of the selected plans. Increasing the number of inference iterations does not have a major impact on mean suboptimality. The median suboptimality has a sharp decline from 5 to 10 inferences but does not show a significant improvement with larger values. The changes in plan runtimes are not significant either. This means the accuracy provided by 10 invocations should be sufficient for selecting robust plans. In addition, this observation demonstrates that the proposed approach is robust to a limited number of predictions. Lastly, Figure 9d demonstrates the inference overheads of the risk-aware strategies compared with the baselines. Overheads for both strategies are less than 6 milliseconds per query. Note that plan selection by suboptimality risk is only slightly more expensive than doing so by the conservative strategy.
@@
Overview of Findings Our experiments demonstrate that Roq produces predictions with a higher level of accuracy and correlations with runtime. They also show that the proposed risk quantification methods and plan selection strategies have significant benefits in selecting robust plans. In addition, they demonstrate the roles of accounting for either plan or estimation risks in risk-aware strategies. Furthermore, the proposed strategies demonstrate significant robustness to shifts in workload characteristics. Lastly, we show that the quality of the resulting plans is robust to a limited number of inference iterations used to quantify risks. Therefore, the proposed methods can be used with minimal inference overheads.

### v3

@@ 5.1.4. Baselines
We compare Roq against a classical RDBMS optimizer, that of IBM Db2, as well as four prominent works, Neo (Marcus et al., 2019), Bao (Marcus et al., 2024), Lero (Zhu et al., 2023), and Balsa (Yang et al., 2022). Neo and Bao are selected due to their demonstrated robustness to errors in input features. Balsa is selected due its to demonstrated robustness to workload shifts. Lero is selected due to its demonstrated minimization of regressions. Each one is evaluated for its ability to select a robust plan from the same set of plans for a given query. To this end, the search space is identical for all approaches and is enumerated as explained.
@@
Neo is a fully learned query optimizer that uses deep neural networks and reinforcement learning to generate efficient query execution plans. It starts by bootstrapping from traditional optimizers and then continuously refines its decision-making by learning from actual query execution times  [^2] Our implementation of Neo uses one-hot encoding for table representation. In order to provide it with the knowledge of the underlying data, we add cardinality estimates to encodings for each plan node..
@@
Balsa is a learned query optimizer that avoids relying on expert demonstrations. It bootstraps from a simple simulator and then fine-tunes its deep reinforcement learning model using real query execution latency feedback.
Lero is a more recent work that proposes a learning-to-rank approach to a learned query optimizer. It learns to compare pairs of plans in the search space and create a total order. The model is trained using a pairwise ranking loss function. Additionally, it enumerates pairs of subplans traversed during search and uses them as additional samples to train the model.
@@ 5.1.5. Evaluation Measures
We design five sets of experiments to evaluate Roq based on prediction error, correlations with runtime, robustness to workload shifts, the inference overheads, and an ablation study on the role of plan and estimation risks.
@@
Correlation. We measure the correlations between the estimates (or the optimizer’s Cost) and the actual execution times using Spearman’s rank coefficient rather than the more widely used Pearson’s coefficient. While Pearson’s coefficient measures the linear relationships between two variables, Spearman’s coefficient measures the monotonic relationships, whether linear or not. Therefore, given the estimated costs or execution times are used to rank the plans, Spearman’s coefficient is a more suitable measure of correlation.
Plan Suboptimality. The quality of plans is evaluated using the suboptimality measure, suggested for end-to-end robustness (Haritsa, 2020). The suboptimality of a plan $p_{i}$ is computed with respect to the best plan $p*$: $\text{Subopt}(p_{i})=\frac{ET(p_{i})}{ET(p^{*})}\in\mathbb{R};1<\text{Subopt}(p_{i})<\infty$, where $ET(p_{i})$ is the execution time of the plan $p_{i}$, and $p*=argmin_{i\in\{1,\ldots,S\}}ET(p_{i})$. The tail-end of the distribution of suboptimality over all plans, selected by a method, characterizes the worst-case performance degradation of that method compared with an oracle that always selects the best plan.
Runtime. Improving the tail-end suboptimality as a measure of robustness may lead to picking plans that are more robust but more expensive than the optimal plan. While the main goal of our work is to improve robustness, we need to demonstrate that robust query optimization does not cause a major slowdown for a given workload. Therefore, the total runtime of the workload is measured for each approach.
Inference Overheads. Using the risk-aware plan evaluation algorithms has an overhead for the inference time since multiple predictions must be made using the base model. We measure the inference time for these algorithms for various inference iterations while evaluating the impact on plan suboptimality and runtime.
@@ 5.2. Experimental Results
Prediction Accuracy. The average predictive performance of Roq’s base model (without uncertainty quantification) consistently outperforms the baselines. Roq’s prediction’s q-error is superior to the baselines in all tested scenarios based on CEB, JOB, and DSB, as shown in Figure 5a.  [^3] Note that Lero is not included in this evaluation as it is a learning-to-rank model, and that, similar to the optimizer’s cost, the absolute values of its predictions do not correspond to latency. Therefore, its q-error is not meaningful. The correlation of the predicted runtime with the actual runtime is more significant for Roq compared with either the baseline ML-based approaches or the Db2 optimizer’s cost (denoted as Cost), as shown in Figure 5b. In these evaluations we only use the raw predictions from each model and do not use MC dropout variational inference to obtain a more accurate prediction.
Figure 5. Roq’s predictive performance vs. the baselines. (a) q-error, and (b) Spearman’s correlation with respect to latency. The Cost baseline refers to the optimizer’s cost estimate
@@
Model Architecture. Better average predictive accuracy may not necessarily lead to selecting a better plan. Therefore, we perform experiments to evaluate the quality of the generated plans based on their suboptimality and runtime. The plans selected by the risk-aware strategies using Roq demonstrate greater robustness based on tail-end plan suboptimality and increased runtime improvements compared with the baselines, as shown in Figures 6a and 6b, respectively.
To isolate the impact of the plan selection strategies as implemented by Algorithms 1, 2, and 3, Roq is tested once without risk quantification, similar to the baselines. This scenario is denoted as Roq-Base. On average, Roq’s base model improves the mean and the $99^{th}$ percentile of suboptimality compared with the optimizer by 26.9% and 28.0%, respectively. In addition, it improves the same measures compared with the ML-based baselines on average by 30.8.3% and 44.9.8% respectively, and it improves runtime compared with the ML-based baselines on average by 22.5%. These improvements are attributed to Roq’s architecture using GNNs for producing node embeddings from the join graph, which enables enhanced generalization to the unseen test queries.
@@
Figure 6. The performance of plans selected by each approach illustrated by (a) suboptimality distribution and (b) improved, regressed, and unchanged percentage of queries, as well as total runtime improvement percentage compared with Db2’s plan
@@
Plan Selection Strategies. The risk-aware plan selection strategies effectively increase the number of improved queries and improve the overall performance and robustness. As demonstrated in Figure 6, in comparison with the ML-based baselines, plan selection by subopt risk (denoted by Roq-Risk) yields the greatest level of robustness by improving the mean and the $99^{th}$ percentile of suboptimality on average by 24.5% and 44.4% respectively. Conservative plan selection (denoted by Roq-Cons) improves the same measures by 22.5% and 44.4% respectively. These two strategies improve runtime compared with the ML-based baseline on average by 23.0% and 17.3% respectively. Roq-Risk and Roq-Cons improve the $99^{th}$ percentile of suboptimality compared with Roq-Base by 19.6% and 19.4%, respectively. While both Roq-Risk and Roq-Cons improve the tail-end performance similarly, the former was more effective at improving the overall runtime. These two strategies are combined with plan pruning based on plan and estimation risks, (denoted by Risk-Roq-Prun and Risk-Cons-Prun). Pruning did not have a significant impact on the performance of either Roq-Risk or Roq-Cons over the three benchmarks. This is expected as these strategies adjust the predictions according to their uncertainties and effectively eliminate risky predictions by default.
@@
Table 3. Comparing the impacts of the risk-based plan selection strategies using model, data, and total uncertainties in suboptimaility distribution for CEB, JOB, and DSB queries.
CEB JOB DSB
Suboptimality median mean 95th 99th median mean 95th 99th median mean 95th 99th
base model 1.05 1.16 1.44 2.68 1.18 1.26 1.86 2.87 6.33 5.85 13.35 16.58
conservative model uncertainty 1.14 1.16 1.59 1.85 1.17 1.20 1.72 1.94 5.49 6.02 12.87 14.75
conservative data uncertainty 1.10 1.13 1.27 1.80 1.17 1.18 1.56 1.83 8.44 7.05 13.94 17.10
conservative total uncertainty 1.10 1.13 1.27 1.80 1.17 1.18 1.56 1.84 4.60 5.10 11.96 14.96
risk model uncertainty 1.14 1.16 1.47 1.85 1.17 1.21 1.72 1.94 4.32 5.72 12.03 14.08
risk data uncertainty 1.09 1.13 1.44 1.80 1.17 1.21 1.72 1.94 6.50 6.60 12.97 15.28
risk total uncertainty 1.14 1.16 1.47 1.85 1.17 1.19 1.64 1.94 3.76 5.36 11.15 14.17
@@
Plan vs. Estimation Risks. To evaluate the roles of the plan risk (quantified by data uncertainty) and estimation risks (quantified by model uncertainty) on plan selection, we run an ablation study in which the risk-aware plan selection strategies are tested using either model, data, or total uncertainties. The quality of the resulting plans is evaluated based on mean, median, 95th and 99th percentiles of the suboptimality of the resulting plans for all three tested benchmarks (Table 3). Using either model or data uncertainty for quantifying risks provides significant improvements compared with the base model and using total uncertainties provides maximum improvements. This suggests that the plan and estimation risks each play a role in improving Roq’s performance and the impact is maximized if they are used together. In addition, the impact of accounting for plan risk is more significant for achieving better runtime performance since it effectively avoids long-running plans and thus significantly reduces regressions.
@@
Robustness to Workload Shifts. A major challenge of using ML-based techniques in query optimization is dealing with the problem of generalization to out-of-distribution test samples. A learned model becomes obsolete if the characteristics of the test query differ significantly from the ones used in training. While model retraining is possible, it takes time, and therefore the optimizer would typically need to still rely on the obsolete model in the meantime. We evaluate the impact of a workload shift for Roq in comparison to the baselines using two simulations.
In a minor workload shift simulation, the baseline model is trained using all query templates in the DSB benchmark. The baseline model is tested on making predictions for two selected templates query025 and query101, due to their unique characteristics 444query025 is a 7-join query with a complex join graph involving three aliases of the ‘DATE_DIM‘ table, each used for fine-grained temporal slicing of a different fact table with overlapping time windows. It includes many-to-many joins between the fact tables STORE_SALES and STORE_RETURNS using multiple join predicates. query101 is a query with 9 joins, involving a self-join over the DATE_DIM table and many-to-many joins between the fact tables STORE_SALES, STORE_RETURNS, and WEB_SALES.. The second model is trained using the other 13 query templates in the DSB and tested on the two selected templates.
@@
Figure 8. JOB runtime changes after a severe workload shift demonstrated by (a) runtime improvements vs. regressions (b) workload cumulative runtime
@@
A major workload shift scenario is simulated using the CEB and JOB. The baseline model is trained and tested on JOB. This represents an ideal scenario where the test workload is identically distributed as the training workload. The model representing a workload shift is trained on CEB and tested on JOB. This constitutes a major workload shift as the query templates in CEB are entirely different than those in JOB. In addition, our CEB queries have up to 12 joins while JOB has up to 28 joins.
The results shown in Figure 7 demonstrate Roq’s robustness to out-of-distribution data caused by a minor workload shift in the first scenario. Figures 7a and 7b show that while the baseline models’ q-error and correlation degrade significantly with a minor workload shift, Roq’s predictive performance degradation is minimal. These results confirm a greater robustness of Roq to out-of-distribution data caused by workload shifts. We also evaluate the impact of workload shift on runtime in the major workload shift scenario. As demonstrated in Figures 8a and 8b, the performance of three out of four baseline models degrades significantly. Lero’s performance slightly improves, which can be attributed to the fact that it does not need to make accurate predictions as long as it can properly rank the plans. Roq’s base model’s performance also improves compared to the baselines. This can be attributed to the use of GNNs for learning generalizable table and query representations from the CEB workload and use that effectively to make accurate predictions on JOB queries and the fact that our CEB workload includes 1000 queries compared to JOB that includes 113. The additional samples enable a greater performance on an unseen test set when generalizable representations are learned. The Conservative and the Risk strategies, further reduce regressions leading to greater net improvements in runtime. Therefore, Roq and its risk-aware strategies meet the design objective (c) of demonstrating robustness based on well-defined measures.
@@
Inference Overheads. Capturing estimation risk using the MC Dropout method requires multiple inferences. A higher number of inferences should provide a more accurate estimate of the expected values and the uncertainties. However, this incurs additional overhead for the query compilation phase, which is on the critical path of query execution and must be kept at a minimum. We experiment with inference iterations ranging from 5 to 100. We use plan selection by suboptimality risk as it is expected to have a larger overhead compared to the conservative plan selection strategy, given the more complex computations. We evaluate the inference time, the runtime, and the suboptimality of selected plans. We report an average over 10 runs for each experiment. Inference times are captured using a single thread on the CPU, typically available for query compilation. As shown in Figure 9a, they grow substantially with the number of iterations. Figures 9b and 9c illustrate the impact of increasing the number of inference iterations on the suboptimality and the runtime of the selected plans. Increasing the number of inferences does not have a major impact on mean suboptimality. The median suboptimality has a sharp decline from 5 to 10 inferences but does not show a significant improvement with larger values. The changes in plan runtimes are not significant either. This means the accuracy provided by 10 inferences should be sufficient for selecting robust plans. This demonstrates that our approach is robust to making a limited number of inferences. Figure 9d demonstrates the inference overheads of the risk-aware strategies compared with the baselines  [^5] Balsa and Lero use Neo and Bao’s value networks respectively. Therefore, their inference overheads are identical to Neo and Bao. While Lero is a learning to rank model, it does not require multiple inferences to create a total order (Zhu et al., 2023).. This time measurement does not include plan generation, as this is identical for all alternatives. Overheads for Roq-Cons and Roq-Risk are around 70 and 89 milliseconds, respectively. As expected, plan selection by suboptimality risk is slightly more expensive than doing so by the conservative strategy. Note that these measurements are performed based on a prototype code in Python and can be significantly reduced with more optimized implementations. Therefore, Roq and its risk-aware strategies meet the design objective (a) of limiting the compilation overheads at a practical level.
Overview of Findings. Roq produces predictions with a higher level of accuracy and correlations with runtime. The risk quantification methods and plan selection strategies have significant benefits in selecting robust plans. Our experiments show the role of accounting for plan or estimation risks in the plan selection strategies. The proposed strategies demonstrate significant robustness to shifts in workload characteristics. The quality of the resulting plans is robust to a limited number of inference iterations used to quantify risks. Therefore, the proposed techniques can be used with minimal inference overheads. Our observations suggest that Roq-Risk and Roq-Cons demonstrate similar results. However, from our experience, Roq-Risk is preferable as it is non-parametric if the computation overheads can be tolerated. Also, pruning by plan or estimation risk does not provide significant improvements for these strategies. Lastly, our ablation study suggests that accounting for Plan Risk is more beneficial than accounting for Estimation Risk.